OpenStudy (anonymous):

Solve the system of equations for x and y by substitution. y = 3x + 6 and y = 5x + 10

OpenStudy (anonymous):

y=3x + 6 y=5x+10 3x+6=5x+10 3x+6-10=5x 3x-4=5x -4=5x-3x -4=2x x=-2 y=3(-2)+6 y=0

OpenStudy (anonymous):

Thanks at @kevin4762